5) (14 points) For the circuit shown below, find the current through R. R - 402 R -202 R-60
Answer #1

Byum R₂=22 Ri=402 . R₂ = 6-2 4V Iov YIs At the junction B, I = I + Iz 0 Applying Kirchoffs voltage law for ABEFA, -IR - Iz R22Iz = 18 E = 1 A substituting Iz= 1 A in @ we get, Now -4 IL-1009) +10 = 0 => -412 = 90 - 10 = 90 - 10 _412 = -20 and I = Iz